OpenStudy (anonymous):

find the indicated partial derivative for each function. fx, fy, fxx, fyy, fxy, fyx a) f(x,y) 3x^4y + 2xy^2- 5xy+8x-9y

OpenStudy (owlfred):

Hoot! You just asked your first question! Hang tight while I find people to answer it for you. You can thank people who give you good answers by clicking the 'Good Answer' button on the right!

OpenStudy (anonymous):

fx= 12x^3y+2y^2-5y+8; fy=3x^4+4xy-5x-9; fxx=\[\delta fx/\delta x\]=36x^2y; fyy=4x; fxy=d(fx)/dy =12x^3+4y-5; fyx=12x^3+4y-5;

OpenStudy (anonymous):

fyx = fxy , mixed partial derivatives

OpenStudy (anonymous):

you get a nice tree when you start doing f xxx , and mixed partial derivatives